OpenStudy (melissa07):

What are the sine, cosine, and tangent of 3 pi over 4 radians?

OpenStudy (melissa07):

@YanaSidlinskiy

OpenStudy (melissa07):

@helpmemath123

OpenStudy (yanasidlinskiy):

4pie/3 = 240 degrees sin 240 = - sqrt 3/2 cos 240 = -1/2 tan 240 = sqrt 3

OpenStudy (melissa07):

thats not any of the choices A. sin Θ = square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = -1 B. sin Θ = negative square root 2 over 2; cos Θ = square root 2 over 2; tan Θ = 1 C. sin Θ = square root 2 over 2; cos Θ = negative square root 2 over 2; tan Θ = 1 D. sin Θ = negative square root 2 over 2; cos Θ = square root 2 over 2; tan Θ = -1

OpenStudy (melissa07):

wouldn't 3pi/4=135 degrees?

OpenStudy (melissa07):

@shinebrightlikeadimon @Lyrae

OpenStudy (melissa07):

@whpalmer4 @bbbbbbbbbbbbbbbbbb

OpenStudy (anonymous):

your flvs? @Melissa07

OpenStudy (melissa07):

yes.

OpenStudy (anonymous):

me too what lesson i can help you @Melissa07

OpenStudy (melissa07):

10.01 Introduction to the Unit Circle

OpenStudy (anonymous):

do you have a kik so it will be easy to help you through?

OpenStudy (melissa07):

No

OpenStudy (anonymous):

oh ok email me through the openstudy email then

OpenStudy (anonymous):

#1 #2 #3 In that order

OpenStudy (whpalmer4):

Okay, do you have a unit circle diagram?

OpenStudy (melissa07):

no i don't @whpalmer4

OpenStudy (melissa07):

oh okay i think i understand now , thank you ! @whpalmer4

OpenStudy (whpalmer4):

Make yourself some flashcards and drill the unit circle every day...